OpenStudy (anonymous):

How do you solve 16^1/4?

OpenStudy (anonymous):

Can someone show me all the steps plz?

OpenStudy (anonymous):

@geerky42

OpenStudy (anonymous):

No

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (lyrae):

\[\large 16^{\frac{ 1 }{ 4 }} = (2^4)^{\frac{ 1 }{ 4 }} = 2^{\frac{ 4 }{ 4 }} = 2\]

geerky42 (geerky42):

That's cruel, old pedo man

OpenStudy (jdoe0001):

\(\large \bf a^{\frac{{\color{blue} n}}{{\color{red} m}}} = \sqrt[{\color{red} m}]{a^{\color{blue} n}} \)

OpenStudy (anonymous):

LMAO @geerky42

OpenStudy (anonymous):

And Thank You @Lyrae just what I needed lol

OpenStudy (lyrae):

Yw :)

OpenStudy (luigi0210):

I was going to say logs, but then I realized it wouldn't work ._.